OpenStudy (anonymous):

8y + 14 = 3x 5x = 4y simultamoes eyuation

OpenStudy (anonymous):

@mathmale @sarahlane235

OpenStudy (anonymous):

x=14/3 and y=0

OpenStudy (anonymous):

Simultamoes eyuation? o_O

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

What do you have to use, substitution or elimination?

OpenStudy (anonymous):

y = 3x - 14/8 is given to be used

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Yes, but what method is it asking you to use?

OpenStudy (anonymous):

substitide

OpenStudy (anonymous):

Ok..so isolate one of the variables and substitute it into the other.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So what have you done so far?

OpenStudy (anonymous):

show me the working and the answer plz

OpenStudy (anonymous):

Show me what you've done so far?

OpenStudy (anonymous):

I'm not just going to do everything for you.

OpenStudy (anonymous):

5x = 4y 5x = 4(3x - 14/8) 5x/4 = 3x - 14/8 8(5x) = 4(3x - 14) 40x = 12x - 56 x = -2 now im stucking i know nothing what to do now

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

There's some problem with the algebra in the equation you used to substitute. Do you see what it is? Since you're using y, what is wrong with what you used?

OpenStudy (anonymous):



OpenStudy (anonymous):

Alright, isolate y, and do it in steps.

OpenStudy (anonymous):

8y= 3x -14 y= 3x - 14/8 now what

OpenStudy (anonymous):

@Jack1

OpenStudy (anonymous):

@wolf1728 @iambatman

OpenStudy (anonymous):